OpenStudy (anonymous):

!?!? How do you even do this without a calculator? square root "a"/ 1+square root "a"

OpenStudy (anonymous):

Here is a better look at it \[\sqrt{a}\div 1+\sqrt{a}\]

OpenStudy (anonymous):

Is it being divided by the 1+sqrt a? or just 1 and later sqrt a is added?

OpenStudy (anonymous):

\[(\sqrt{a})/(1+\sqrt{a})\]

OpenStudy (anonymous):

then ur in luck! The sqrt a's cancel out, and ur left with 1/1, aka1

OpenStudy (anonymous):

u see what I did?

OpenStudy (anonymous):

I see what you did, but my book gives the answer \[(\sqrt{a}-a)/ 1-a\]

OpenStudy (anonymous):

thats weird. Dunno what they did